Please help me on this problem: Solve the linear system using the substitution method.
x=16-4y
3x+4y=8

Answer by boilpoil(127)

x=16-4y
3x+4y=8 ---Since x is already the subject, we'll substitute x into second equation

3(16-4y)+4y=8
48-12y+4y=8
-8y=-40
y=5

x=16-4(5)
x=16-20
x=-4